### v1

@@ Abstract
Language processing is at the heart of current developments in artificial intelligence, and quantum computers are becoming available at the same time. This has led to great interest in quantum natural language processing, and several early proposals and experiments. This paper surveys the state of this area, showing how NLP-related techniques including word embeddings, sequential models, attention, and grammatical parsing have been used in quantum language processing. We introduce a new quantum design for the basic task of text encoding (representing a string of characters in memory), which has not been addressed in detail before.
As well as motivating new technologies, quantum theory has made key contributions to the challenging questions of “What is uncertainty?” and “What is intelligence?” As these questions are taking on fresh urgency with artificial systems, the paper also considers some of the ways facts are conceptualized and presented in language. In particular, we argue that the problem of ‘hallucinations’ arises through a basic misunderstanding: language expresses any number of plausible hypotheses, only a few of which become actual, a distinction that is ignored in classical mechanics, but present (albeit confusing) in quantum mechanics.
@@
In early 2024, quantum computing and AI are two of the most rapidly-moving and talked-about areas of science and technology.
The availability of dialog systems based on large language models (LLMs) has raised the profile of natural language processing (NLP) to a new historic high, developing and expanding very quickly. This expansion has led to AI models being deployed as systems and introduced as components in new ways, leading to improvements and efficiencies, but also mistakes and concerns. Thus the demand for improvements in AI is at an all-time high, with a renewed focus on reliability and trust.
Quantum theory offers new forms of mathematical modeling, computation and communication. Mathematical models for language operations motivated explicitly by quantum theory have been used in information retrieval (Van Rijsbergen, 2004; Sordoni et al., 2013), logic and disambiguation (Widdows and Peters, 2003), and language composition (Clark and Pulman, 2007; Coecke et al., 2010). Similar models have been developed in many social sciences and demonstrated successful results over classical alternatives, long before any such models were implemented and run on quantum computers. More abstractly, entire classes of quantum machine learning (ML) models have been theoretically shown to have more expressive power than comparable classical models (Coyle et al., 2020; Yu et al., 2023). Running basic NLP algorithms on quantum has become possible only in the last few years, with early-stage results reported by Lorenz et al. (2021); Widdows et al. (2022).
This paper is intended as an introduction to this landscape, for those interested in language processing and quantum computing, but not necessarily specialists in either. The paper is in three main sections. Firstly, we give a brief introduction to quantum gates and circuits, and a detailed example of how they might be used for representing the sequences of characters that make natural language texts. This gives a glimpse of some of the wonder, and some of the challenges, of quantum computing. The main body of the paper surveys ways in which other aspects of language processing have already been modeled on quantum computers, including embedding vectors, sequences, attention, and grammatical structure. This gives a snapshot of of where quantum NLP has got to at this stage of the NISQ era. Finally, the paper discusses the challenges of choosing and distinguishing between the hypothetical and the actual. This has taken on fresh urgency in AI systems for fact-checking, to avoid mistaking ‘hallucinations’ for assertions. We note that language models are designed to produce both hypothetical and actual statements, and that quantum mechanics is a better starting point than classical mechanics for modeling this.
2 Quantum Basics in the NISQ Era
@@
In mathematical terms, the key features that distinguish quantum from classical computers are superposition and entanglement. Superposition can be realized in a single qubit: the state is a superposition of the states and , where and are complex numbers, with . Each single-qubit logic is a linear operator that preserves the orthogonality of the basis states and this normalization condition, and the group of such operators is , the group of complex unitary matrices. Single-qubit gates that feature prominently in this paper are shown in Figure 2. So single-qubit gates coherently manipulate the superposition state of an individual qubit.
@@
A quantum circuit consists of a register of qubits, and a sequence of logic gates that act on these qubits. Some of the basic gates used in this paper are shown in Figures 2 and 2.
The Pauli- gate is commonly used to flip a qubit between the and gates, which is why it is also sometimes called the quantum NOT gate. -gates directed at different qubits can be used to prepare an input state representing a binary-valued vector: the state is prepared by applying an -gate to each of the qubits to be switched to the state.
@@
Other probabilities, anywhere in the range , can be arranged by using fractional rotations (which in practice might involve sending just the same laser-pulse instructions but for different carefully-tuned time periods). An example is given in the gate. Several variational quantum algorithms work by gradually optimizing such parameters.
@@
The Pauli-, Hadamard, and CNOT gates are self-inverse: performing these operations twice gives the identity map.
@@
By assembling several 1- and 2-qubit gates, more qubits can become entangled, and we can define multi-controlled gates. For example, the 3-qubit Toffoli gate in Figure 2 flips the lowest qubit if the top qubit is in the state and the middle qubit in the state. It is relatively easy mathematically to start arranging such gate recipes into higher-level operations: for example, the 3-qubit gate acts as a logical AND, from which simple binary arithmetic can be constructed. However, developers must be careful when using such constructions, because gate complexity and errors can easily build up. In practice, it takes 5 CNOT gates and 9 single-qubit gates to assemble the 3-qubit Toffoli gate, so if gate errors are above 1%, the error-rate in a circuit with just 3 Toffoli gates would be over 50%. In 2024, gate error rates tend to be much better than this, but still, circuits in the NISQ era today do not reliably run more than a few hundred gates. Fault-tolerant gates and quantum error correction will eventually improve this situation, but at the moment, quantum circuit development tends to tradeoff between sophistication (more gates introduces more tunable parameters) and reliability (fewer gates gives fewer errors).
@@ 3 A Quantum String Encoding Example
Character and string encoding is one of the most basic tasks in language processing, and this section gives a worked example of how this might be performed on a quantum computer using some of the standard quantum circuit and gate patterns introduced in the previous section. This makes a good case-study of some of the promise and challenges of quantum computing, and (as far as we know) is the first such proposal for representing text strings in a quantum computer (comparable to the use of ASCII or Unicode specifications in mainstream classical computing).
Other methods of encoding the meanings of texts in quantum NLP work have been devised, such as vector embeddings for use in machine learning classifiers, and several such techniques will be surveyed in later sections. The example quantum circuit designs in this section are for the (much older) protocol of representing words as a sequence of characters chosen from a relatively small character set. Some of the quantum word-encoding models based on a sequences and embeddings will be surveyed in later sections of this paper.
@@
One data structure we can use as a building block in quantum circuits is binary positional notation for integers. For example, in a -qubit register, the state could represent the decimal number 10 (if read from the left) or 5 (if read from the right). This convention was used right at the beginning of quantum computing, in Feynman’s proposal of how to build a quantum adder circuit (Feynman, 1985), and is part of Shor’s integer factoring algorithm (Shor, 1994). These numbers can be represented in quantum circuits using -gate ‘bit-flip’ operations on the corresponding qubits, as in Figure 3.
@@
Our string encoding design works by entangling a “position” register that records where a character appears, along with an “alphabet” register that says which character appears in that position. So instead of a sequence , the string is represented as a tensor product , where is the state representing the position, and is the state representing the character. A similar pattern is used by Amankwah et al. (2022) for representing images, and the encoding is called QPIXL. QPIXL uses one register to specify the location of a pixel in an image, entangled with another register saying which ‘channel’ (e.g., red, green, or blue) is being referred to. So, QPIXL also keeps ‘what it is’ and ‘where it is’ in separate registers and entangles these. Emphasizing this similarity, we call our string encoding protocol QPOSTR, pronounced ‘Q-poster’, meaning “Quantum Positional String”.
The implementation of this formula as a quantum circuit component for encoding the string cab is outlined in Figure 5. The top 2 qubits form the “position” register, and the values on the control qubits are represented by the open and closed circles, , . Starting with the top as the least significant bit, the control states are , , , etc. The bottom 2 qubits form the ‘character’ or ‘alphabet’ register. Each letter in the alphabet is mapped to a number corresponding to its position, so the gate-recipe for each character is like one of the simple circuits shown in Figure 3. To encode the string cab, a character register of 2 qubits suffices. To encode 26 letters, a character register of 5 qubits would be required (since ), and for the ASCII character set, 7 qubits would be needed.
If the circuit above is prepared in the conventional state, the first gate controlled on the state is the only one active, and the circuit output will be 00 (in the position register) and 11 (in the character register), saying just “the zeroth character was a c.” To prepare a superposition of the characters in all of the positions, the position register is prepared in a uniform distribution over all the available character positions, using a standard array of Hadamard (H) gates. This gives the full circuit for encoding the string cab in Figure 6. Character positions beyond the length of the string are untouched, or left with character ‘0’ in that position. Using the convention that character ‘0’ represents a space, this is equivalent to padding a string with trailing zeros to make its length a power of 2.
@@
As a thought experiment, we can use GPT-3 training metadata to demonstrate this savings. The size of the training dataset is reported at 300B tokens (Brown et al., 2020), so a generous estimate of 12 characters-per-token allows for character-positions, for which the positional encoding fits in 42 qubits. There are currently 149,813 Unicode characters, so even this alphabet fits in 18 qubits, which means that the entire 45TB training dataset of GPT-3 can fit into a mere qubits!
@@
If this process could be repeated many times, and if the output qubits can be reset to independently of the other qubits in the register, eventually the entire input string or any parts of it can be read out to classical memory. Superficially, this extension of the readout circuit gives a circuit that demonstrates that the whole original string can be recovered from the QPOSTR quantum encoding. In other words, it looks superficially as if we can recover an exponentially-long string ( characters) from an encoding that uses qubits!
@@
Thus, QPOSTR gives an exponential space advantage over the classical alternative. However, it still takes a length of time at least linear in the length of the string to run the quantum circuit that “loads” the string into quantum memory. Like many quantum encodings, QPOSTR only gives an exponential space saving, which offers no obvious practical advantage without a corresponding time saving. Thus, even if we were to encode the GPT-3 training data, it is unclear at this moment what savings could come from doing so. That said, we optimistically speculate that future algorithms or data structures may be able to better take advantage of this encoding.
More generally, the challenge of preparing a quantum memory that can be maintained and successively queried is sometimes described as research in QROM and QRAM (Giovannetti et al., 2008; Babbush et al., 2018). Minimizing the number of gate operations is a key goal in such work, and the position encoding used in QPOSTR can be regarded as one of the ‘simple (but suboptimal)’ encoding methods described in Fig 3 of Babbush et al. (2018). The task they are interested in is encoding electronic spectra, but they also consider encoding ‘words’ as an example toy problem. The extra step that QPOSTR takes is explicitly to map register values to alphabetic characters, which enables such a unitary positional encoding to represent a text as a sequence of alphabetic characters.
Rather than demonstrating a new ‘quantum advantage’, the QPOSTR example is intended to showcase some of the excitement, but also some of the gotchas of quantum computing. It is astonishing and exciting that an exponentially-long string can be encoded like this at all, but once the engineering caveats around that statement are properly understood, we see that the explicit information we can recover from this representation is much smaller. The following sections consider other representations for text, including vectors of coordinates (‘embeddings’), sequences, and grammatical trees.
@@ 4 Word Embeddings and Text Classification
Representing words as vectors of coordinates is a technique that goes back at least to the 1960s and early information retrieval systems (Salton and McGill, 1983). Semantic properties of vectors in lower-dimensional projections were analyzed in the 1990s (Landauer and Dumais, 1997), and by the early 2000s, overlaps between the logic of word vectors in information retrieval and state vectors in quantum mechanics had been explicitly recognized (Van Rijsbergen, 2004; Widdows, 2004). In the past decade, embeddings for classical NLP have jumped from having a resurgence in academia to becoming massively mainstream in industry (Mikolov et al., 2013a; Bridgwater, 2023; Metinko, 2023). Naturally, this suggests embeddings could be just as central for QNLP, especially since the mathematics of vectors and tensors has become a common language for both AI and quantum computing (Widdows et al., 2021).
@@
Word2ket (Panahi et al., 2019) was introduced as a quantum-inspired solution for compressing embeddings. A ‘tensor network’ is a decomposition of a high-dimensional tensor into an approximate product of lower-dimensional tensors or vectors. For example, if then can be represented using approximately coordinates, rather than . (More precisely, in matrix coordinates, for column vectors and corresponding to appropriately reshaped and , and is a reshaped version of ; see Hitchcock (1927); Van Loan (2000) for details.) Word2ket uses tensor networks to create low-dimensional approximations for individual word vectors, and entire vocabularies. As an aside, this mathematical initiative continues: for example, Tomut et al. (2024) report using tensor networks to compress the parameters of an LLM (LlaMA-2 7B model) to 30% of its original size while retaining over 90% of the original accuracy.
These methods are quantum-inspired, in the sense of drawing deliberately on quantum mathematical models, but running on classical computers. Next we discuss in further detail techniques relating to embeddings intended for use on actual quantum devices, treating separately the topics of building quantum embeddings and using them.

### v2

@@ Abstract
Language processing is at the heart of current developments in artificial intelligence, and quantum computers are becoming available at the same time. This has led to great interest in quantum natural language processing, and several early proposals and experiments.
This paper surveys the state of this area, showing how NLP-related techniques have been used in quantum language processing. We examine the art of word embeddings and sequential models, proposing some avenues for future investigation and discussing the tradeoffs present in these directions. We also highlight some recent methods to compute attention in transformer models, and perform grammatical parsing. We also introduce a new quantum design for the basic task of text encoding (representing a string of characters in memory), which has not been addressed in detail before.
Quantum theory has contributed toward quantifying uncertainty and explaining “What is intelligence?” In this context, we argue that “hallucinations” in modern artificial intelligence systems are a misunderstanding of the way facts are conceptualized: language can express many plausible hypotheses, of which only a few become actual.
@@
In early 2024, quantum computing and AI are two of the most rapidly-moving and talked-about areas of science and technology. The availability of dialog systems based on large language models (LLMs) has raised the profile of natural language processing (NLP) to a historic high, developing and expanding very quickly. This expansion has led to AI models being deployed as systems and introduced as components in new ways, leading to improvements and efficiencies, but also mistakes and concerns. Thus the demand for improvements in AI is at an all-time high, with a renewed focus on reliability and trust.
Quantum theory offers new forms of mathematical modeling, computation and communication. Mathematical models for language operations motivated explicitly by quantum theory have been used in information retrieval (Van Rijsbergen, 2004; Sordoni et al., 2013), logic and disambiguation (Widdows and Peters, 2003), and language composition (Clark and Pulman, 2007; Coecke et al., 2010). Similar models have been developed in many social sciences and demonstrated successful results over classical alternatives, long before any such models were implemented and run on quantum computers. More abstractly, entire classes of quantum machine learning (ML) models have been theoretically shown to have more expressive power than comparable classical models (Coyle et al., 2020; Yu et al., 2023), though this does not guarantee improvements in results more generally (Bowles et al., 2024). Running basic NLP algorithms on quantum has become possible only in the last few years, with early-stage results reported by Lorenz et al. (2023); Widdows et al. (2024).
This paper is intended as an introduction to this landscape, for those interested in language processing and quantum computing, but not necessarily specialists in either. Firstly, Section 2 gives a brief introduction to quantum gates and circuits. Section 3 continues with an idealized example of how quantum gates could be used to represent a text string of exponential length in a register of qubits, including some caveats and pitfalls. This gives a glimpse of some of the wonder, and some of the challenges, of quantum computing. The main body of the paper surveys ways in which other aspects of language processing have already been modeled on quantum computers, including embedding vectors, sequences, attention, and grammatical structure (Sections 4–7). This gives a snapshot of of where quantum NLP has got to at this stage of the NISQ era. Finally, Section 8 discusses the challenges of choosing and distinguishing between the hypothetical and the actual. This has taken on fresh urgency in AI systems for fact-checking, to avoid mistaking so-called hallucinations for assertions. We note that language models are designed to produce both hypothetical and actual statements, and that quantum mechanics is a better starting point than classical mechanics for modeling this.
2 Quantum Computing Basics
@@
In mathematical terms, the key features that distinguish quantum from classical computers are superposition and entanglement. Superposition can be realized in a single qubit: the state is a superposition of the states and , where and are complex numbers, with . Each single-qubit logic gate is a linear operator that preserves the orthogonality of the basis states and this normalization condition, and the group of such operators is , the group of complex unitary matrices. Single-qubit gates that feature prominently in this paper are shown in Figure 2. So single-qubit gates coherently manipulate the superposition state of an individual qubit.
@@
A quantum circuit consists of a register of qubits, and a sequence of logic gates that act on these qubits. Some of the basic gates used in this paper are shown in Figures 2 and 2. The Pauli- gate is commonly used to flip a qubit between the and gates, which is why it is also sometimes called the quantum NOT gate. -gates applied to different qubits can be used to prepare an input state representing a binary-valued vector: the state is prepared by applying an -gate to each of the qubits to be switched to the state.
@@
Other probabilities, anywhere in the range , can be arranged by using fractional rotations (which might involve sending just the same laser-pulse instructions, but for different time periods). An example is given in the gate. Several variational quantum algorithms work by gradually optimizing such parameters.
@@
By assembling several 1- and 2-qubit gates, more qubits can become entangled, and we can define multi-controlled gates. For example, the 3-qubit Toffoli gate in Figure 2 flips the lowest qubit if the top qubit is in the state and the middle qubit in the state. It is relatively easy mathematically to start arranging such gate recipes into higher-level operations: for example, the 3-qubit gate acts as a logical AND, from which simple binary arithmetic can be constructed.
However, developers must be careful when using such constructions, because gate complexity and errors can easily build up. In practice, it takes 5 CNOT gates and 9 single-qubit gates to assemble the 3-qubit Toffoli gate, so if gate errors are above 1%, the error-rate in a circuit with just 3 Toffoli gates would be over 50%. In 2024, gate error rates tend to be much better than this, but still, typical circuits today do not reliably run more than a few hundred gates. NISQ-era quantum circuit development tends to tradeoff between sophistication (more gates introduces more tunable parameters) and reliability (fewer gates gives fewer errors). It also leads to designs where classical components are relied upon for many parts of an NLP pipeline, such as storing weights and comparing scores (Widdows et al., 2024).
In quantum machine learning, variational circuits, or parametrized quantum circuits (PQCs), are a particularly clear example of such tradeoffs (Schuld and Petruccione, 2021, Ch 5). Variational circuit designs use i. a quantum circuit with parameters , often implemented as variable gate angles, that can be optimized according to a given loss function; and ii. a classical optimizer, responsible for evaluating the measurement outputs of the quantum circuit, and proposing updates to the parameters . When both classical and quantum components play such a prominent role, the combination is sometimes called a hybrid system or hybrid workflow.
The transition from NISQ to fault-tolerant quantum computing will be gradual, and arguably is already underway: recent months have seen encouraging progress in quantum error correction and memory fidelity (Bravyi et al., 2024). This raises long-term expectations that quantum computers will optimize crucial matrix operations, such as solving systems of equations (Harrow et al., 2009), and quantum singular value transformation (Gilyén et al., 2019). However, it is important to remember that even fault-tolerant quantum computing will come with serious caveats, and in particular, quantum components bring no advantage if their I/O costs outweigh their computational gains (Aaronson, 2015).
Through the rest of this paper, we highlight examples of some of these considerations and design differences as they appear.
@@ 3 A Quantum String Encoding Example
Character and string encoding is one of the most basic tasks in language processing, and this section gives a worked example of how this might be performed on a quantum computer using some of the standard quantum circuit and gate patterns introduced in the previous section. This makes a good case-study of some of the promise and challenges of quantum computing, and (as far as we know) is the first such proposal for representing text strings in a quantum computer, comparable to the use of ASCII or Unicode specifications in mainstream classical computing.
To encode a text of meaningful length, this encoding would require many layers of 2-qubit gates, and this design would require error correction, rather than being a NISQ-era proposal. Other methods of encoding the meanings of texts in quantum NLP work have been devised, such as vector embeddings for use in machine learning classifiers, and several such techniques will be surveyed in later sections. The example quantum circuit designs in this section are for the (much older) protocol of representing words as a sequence of characters chosen from a relatively small character set. Some of the quantum word-encoding models based on a sequences and embeddings will be surveyed in later sections of this paper.
@@
One data structure we can use as a building block in quantum circuits is binary positional notation for integers. For example, in a -qubit register, the state could represent the decimal number 10 (if read left-to-right) or 5 (if read right-to-left). This convention was used right at the beginning of quantum computing, in Feynman’s proposal of how to build a quantum adder circuit (Feynman, 1985), and is part of Shor’s integer factoring algorithm (Shor, 1994). These numbers can be represented in quantum circuits using -gate bit-flip operations on the corresponding qubits, as in Figure 3.
@@
Our string encoding design works by entangling a “position” register that records where a character appears, along with an “alphabet” register that says which character appears in that position. So instead of a sequence , the string is represented as a tensor product , where is the state representing the position, and is the state representing the character. A similar pattern is used by Amankwah et al. (2022) for representing images, and the encoding is called QPIXL. QPIXL uses one register to specify the location of a pixel in an image, entangled with another register saying which channel (e.g., red, green, or blue) is being referred to. So, QPIXL also keeps “what it is” and “where it is” in separate registers and entangles these. Emphasizing this similarity, we call our string encoding protocol QPOSTR, pronounced “Q-poster”, meaning “Quantum Positional String”.
The implementation of this formula as a quantum circuit component for encoding the string cab is outlined in Figure 5. The top 2 qubits form the “position” register, and the values on the control qubits are represented by the open and closed circles, , . Starting with the top as the least significant bit, the control states are , , , etc. The bottom 2 qubits form the “character” or “alphabet” register. Each letter in the alphabet is mapped to a number corresponding to its position, so the gate-recipe for each character is like one of the simple circuits shown in Figure 3. To encode the string cab, a character register of 2 qubits suffices. To encode 26 letters, a character register of 5 qubits would be required (since ), and for the ASCII character set, 7 qubits would be needed.
If the circuit above is prepared in the conventional state, the first gate controlled on the state is the only one active, and the circuit output will be 00 (in the position register) and 11 (in the character register), saying just “the zeroth character was a c.” To prepare a superposition of the characters in all of the positions, the position register is prepared in a uniform distribution over all the available character positions, using a standard array of Hadamard (H) gates. This gives the full circuit for encoding the string cab in Figure 6. Character positions beyond the length of the string are untouched, or left with character “0” in that position. Using the convention that character “0” represents a space, this is equivalent to padding a string with trailing zeros to make its length a power of 2.
@@
As a thought experiment, we can use GPT-3 training metadata to demonstrate this savings. The size of the training dataset is reported at 300B tokens (Brown et al., 2020), so a generous estimate of 12 characters-per-token allows for character-positions, for which the positional encoding fits in 42 qubits. There are currently 149,813 Unicode characters, so even this alphabet fits in 18 qubits, which means that the entire 45TB training dataset of GPT-3 could fit into a mere qubits!
@@
If this process could be repeated many times, and if the output qubits could be reset to independently of the other qubits in the register, eventually the entire input string or any parts of it could be read out to classical memory. Superficially, this extension of the readout circuit gives a circuit that demonstrates that the whole original string can be recovered from the QPOSTR quantum encoding. In other words, it looks as if we can recover an exponentially-long string ( characters) from an encoding that uses qubits!
@@
Thus, QPOSTR gives an exponential space advantage over the classical alternative. However, it still takes a length of time at least linear in the length of the string to run the quantum circuit that loads the string into quantum memory. Like many quantum encodings, QPOSTR only gives an exponential space saving, which offers no obvious practical advantage without a corresponding time saving. Thus, even if we were to encode the GPT-3 training data, it is unclear at this moment what savings could come from doing so. That said, we optimistically speculate that future algorithms or data structures may be able to better take advantage of this encoding.
More generally, the challenge of preparing a quantum memory that can be maintained and successively queried is sometimes described as research in QROM and QRAM (Giovannetti et al., 2008; Babbush et al., 2018). Minimizing the number of gate operations is a key goal in such work, and the position encoding used in QPOSTR can be regarded as one of the “simple (but suboptimal)” encoding methods described in Fig 3 of Babbush et al. (2018). The task they are interested in is encoding electronic spectra, but they also consider encoding “words” as an example toy problem. The extra step that QPOSTR takes is explicitly to map register values to alphabetic characters, which enables such a unitary positional encoding to represent a text as a sequence of alphabetic characters.
Rather than demonstrating a new quantum advantage, the QPOSTR example is intended to showcase some of the excitement, but also some of the gotchas of quantum computing. It is astonishing that an exponentially long string can be encoded like this at all, but once the engineering caveats around that statement are properly understood, we see that the explicit information we can recover from this representation is much smaller.
@@ 4 Word Embeddings and Text Classification
Representing words as vectors of coordinates is a technique that goes back at least to the 1960s and early information retrieval systems (Salton and McGill, 1983). The key theoretical motivation behind such distributional semantics methods is that words that appear in similar contexts tend to have similar meanings Wittgenstein (1953); Firth (1957). Based on their distribution in text, embedding techniques map words to vector spaces, where their similarity is typically measured by the inner product of their corresponding vectors.
Semantic properties of vectors in lower-dimensional projections were analyzed in the 1990s (Landauer and Dumais, 1997), and by the early 2000s, overlaps between the logic of word vectors in information retrieval and state vectors in quantum mechanics had been explicitly recognized (Van Rijsbergen, 2004; Widdows, 2004). In the past decade, embeddings for classical NLP have jumped from having a resurgence in academia to becoming massively mainstream in industry (Mikolov et al., 2013a; Bridgwater, 2023; Metinko, 2023). Naturally, this suggests embeddings could be just as central for QNLP, especially since the mathematics of vectors and tensors has become a common language for both AI and quantum computing (Widdows et al., 2021).
@@
Word2ket (Panahi et al., 2019) was introduced as a quantum-inspired solution for compressing embeddings. A tensor network is a decomposition of a high-dimensional tensor into an approximate product of lower-dimensional tensors or vectors. For example, if then can be represented using approximately coordinates, rather than . (More precisely, in matrix coordinates, for column vectors and corresponding to appropriately reshaped and , and is a reshaped version of ; see Hitchcock (1927); Van Loan (2000) for details.) Word2ket uses tensor networks to create low-dimensional approximations for individual word vectors, and entire vocabularies. This mathematical initiative continues: for example, Tomut et al. (2024) report using tensor networks to compress the parameters of an LLM (LlaMA-2 7B model) to 30% of its original size while retaining over 90% of the original accuracy.
These methods are quantum-inspired, in the sense of drawing deliberately on quantum mathematical models, but running on classical computers.